Fermi National Laboratory
 

Accelerator Update for January 12, 2007 - January 15, 2007

 

Friday January 12

The day shift began with Operations monitoring store 5174 (46.77E30) and a Recycler stash (248.6E10), with the Antiproton Source (Pbar) stacking (51.0E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off due to absorber work, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

MI experts began a study period around 9:47 AM. These studies caused RF trips and injection kicker trips.

Tevatron (TeV) experts began a Tevatron Electron Lens (TEL2) study 11:03 AM.

At 12:18 PM, during a Recycler transfer, Operations reported that a MI RF station (MIRF14) was tripping off. After the stash had finished an expert replaced the station’s cavity tuning controller box.

Operations terminated store 5174 at 1:03 PM.

A TeV RF expert investigated a TeV RF station (TRF8) that had tripped off on Thursday. The expert reported that he couldn’t find anything wrong with this RF station. He replaced the control chassis, because it was a possible cause for the trip.

Operations put the TeV into shot setup at 1:37 PM.

Operations established store 5175 at 3:24 PM with an initial luminosity of 258.23E30. This store set a Record Luminosity.

Operations reported at 3:42 PM that the beam permits and run conditions were in place so that Switchyard (SY) experts could run beam to the Meson absorber in M01 this weekend.

The Recycler completed its last of three stashes for the day shift at 3:40 PM.

The MI studies ended at 4:26 PM.

At 5:37 PM, it was reported that the NuMI horn wasn’t pulsing properly. Operators turned off the NuMI beam and contacted an expert.

Operations turned off the NuMI horn at 6:03 PM per an expert’s request. The experiment continues to take beam in its horn off mode An expert was called in.

A NuMI horn expert replaced a trigger sync module and the horn resumed pulsing normally at 11:32 PM. NuMI resumed taking beam.

At 11:52 PM, the Fermi earthquake monitor, also known as the Tevatron, recorded a 7.9 magnitude earthquake that had occurred about 300 miles south, southwest of the Kuril Islands in Russia. The store wasn’t disturbed.

Saturday January 13

The midnight shift began with Operations monitoring store 5175 (93.72E10) and the stash (172.2E10), with Pbar stacking (8.4E10), and with NuMI taking beam.

At 12:09 AM, a Linac Klystron RF station (KRF6) tripped off and wouldn’t reset. Operators responded and tweaked the pressure regulator, per an expert suggestion, but it didn’t help. Operations called in experts.

At 3:16 AM, the Pelletron tripped off due to vacuum pump trips. Everything reset and turned on, but tripped back off while trying to establish an electron beam. Operators contacted an expert.

Linac experts managed to bring KRF6 back on line around 3:36 AM. NuMI resumed taking beam.

At 3:36 AM, Operations reported that a Pbar lithium lens interlock trip was holding off beam. Operators investigated and found that the real problem was due to a power supply (HV100) interlock problem. Operations called in a power supply expert.

A Pbar electron cooling expert suggested at 5:14 AM that operators restore a data file. This restored the electron-cooling beam.

At 9:35 AM, Operations reported that HV100 tripped off on a SCR fault. An expert investigated.

At 10:05 AM, the NuMI beam permit tripped off for the first of two times in 26 minutes due to a beam position problem. It reset okay both times.

At 10:46 AM, a power supply expert reported a LCW leak in HV100. A hose clamp had come loose. The expert tightened the hose and began drying the power supply.

Operations turned on HV100 at 12:36 PM. Pbar resumed stacking.

TeV experts began a TEL2 end-of-store study at 3:34 PM.

Operations terminated store 5175 at 4:04 PM.

The Recycler completed its second and last stash for the day shift at 4:10 PM.

Operations put the TeV into shot setup at 4:20 PM.

Operations established store 5176 at 6:30 PM with an initial luminosity of 223.53E30.

The Recycler completed its second and last stash for the evening shift at 10:13 PM.

At 10:15 PM, NuMI tripped off for the first of three times in about an hour and 15 minutes due to beam position problems. It reset each time.

Sunday January 14

The midnight shift began with Operations monitoring store 5176 and the stash, with Pbar stacking, and with NuMI taking beam.

Operations reported at 2:05 AM that KRF6 had alarmed on a high temperature alarm earlier on the shift. An expert said that it would be okay as long as the station temperature stayed below 120˚F.

At 2:16 AM, the Pbar electron cooling suffered a vacuum burst. Operations called in an expert.

An electron-cooling expert reported at 5:44 AM that the Pelletron was back on line. He said that he would need to condition the top part of the Pelletron some time after the next store gets established.

The Recycler completed its second and last stash for the midnight shift at 7 AM.

At 7:46 AM, the NuMI horn tripped off. It reset okay.

At 12:25 PM, the NuMI horn tripped off. It reset okay.

At 2:43 PM, Pbar Debuncher cooling problems reappeared. Operations called in experts.

The Recycler completed its third and last stash for the day shift at 4:34 PM.

Operations terminated store 5176 at 4:38 PM.

Operations put the TeV into shot setup at 4:54 PM.

Operations established store 5177 at 6:41 PM with an initial luminosity of 250.32E30.

A Pelletron expert and operator access MI-31 at 8:29 PM to prepare for the Pelletron conditioning.

Operations reported at 8:33 PM that SY experts had established beam to SY absorber.

SY experts established beam to the M01 absorber at 10:12 PM and then stopped tuning for the night.

At 10:18 PM, Operations reported that the TeV sector E2 wet engine was slowing. Operators investigated and found a worn belt and a wobbly flywheel. Operators informed the Cryo Coordinator and the cryo system technicians at CHL.

At 11:57 PM, Pbar experienced more Debuncher stochastic cooling problems. Operations contacted an expert.

Monday January 15

The midnight shift began with Operations monitoring store 5177 and the stash, with Pbar stacking, and with NuMI taking beam.

At 12:59 AM, TRF1 tripped off.

At 1:40 AM, several TeV vacuum pumps tripped at sector B1. They all reset.

At 3:13 AM, KRF6 tripped off on a temperature alarm. It reset okay.

At 3:24 AM, Operations reported that the electron-cooling front-end wasn’t responding. A reboot brought it back.

At 3:40 AM, MI suffered a vacuum burst. It had recovered within three minutes.

At 4:47 AM, Operations received a NuMI horn LCW expansion tank low-level alarm. Operations contacted an expert.

At 5:04 AM, KRF4 tripped off. It took operators about 15 minutes to reboot the VME, recycle the contactor, get the gradient back to normal, and turn the station on.

Operation reestablished beam to M01 at 5:33 AM in order to document the beam line and make a save file. This took about 30 minutes.

At 6:14 AM, the Recycler’s horizontal stochastic cooling amplifier (H4A23) tripped off. Operators investigated and found that the signal suppression measurement program had hung up. The program resumed normal running.

At 6:55 AM, Operations reported that TeV luminosity had taken a significant drop over the last hour and three quarters. Operator investigation hadn’t determined the problem.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/16/2007    email Fermilab